OpenStudy (anonymous):

Solve for x: 5x + 2 = 4x - 9. 7 -7 11 -11

OpenStudy (shamil98):

5x + 2 = 4x - 9 -4x -2 -4x -2 x = -11

OpenStudy (anonymous):

Solve for x: 3(x + 1) = -2(x - 1) + 6. 1 4 5 25